### v3

# Conquering the Communication Constraints to Enable Large Pre-Trained Models in Federated Learning

Abstract
Federated learning (FL) has emerged as a promising paradigm for enabling the collaborative training of models without centralized access to the raw data on local devices. In the typical FL paradigm (e.g., FedAvg), model weights are sent to and from the server each round to participating clients. Recently, the use of small pre-trained models has been shown to be effective in federated learning optimization and improving convergence. However, recent state-of-the-art pre-trained models are getting more capable but also have more parameters. In conventional FL, sharing the enormous model weights can quickly put a massive communication burden on the system, especially if more capable models are employed. Can we find a solution to enable those strong and readily available pre-trained models in FL to achieve excellent performance while simultaneously reducing the communication burden? To this end, we investigate the use of parameter-efficient fine-tuning in federated learning and thus introduce a new framework: FedPEFT. Specifically, we systemically evaluate the performance of FedPEFT across a variety of client stability, data distribution, and differential privacy settings. By only locally tuning and globally sharing a small portion of the model weights, significant reductions in the total communication overhead can be achieved while maintaining competitive or even better performance in a wide range of federated learning scenarios, providing insight into a new paradigm for practical and effective federated systems. Code is available at https://github.com/imguangyu/FedPEFT.
Index Terms:
Federated learning, Parameter-efficient fine-tuning, Vision transformers, Image Classification, Action Recognition.
Figure 1: Process in a federated learning communication round with $\mathbf{M}$ participating clients. We use ViT-Base as an instance to analyze the communication costs. (a) Conventional federated learning framework, where the entire model will be sent during the communication. (b) FedPEFT, which is our proposed parameter-efficient framework for federated learning. Instead of uploading and downloading the entire model during the communication, only a small subset of the model will be updated and communicated, leading to a huge relief in communication, specifically from about 328MB per client per round to about 0.68MB per client per round.
I Introduction
Federated learning (FL) [1] has become increasingly prevalent in the research community, having the goal of enabling collaborative training with a network of clients without needing to share any private data. One key challenge for this training paradigm is overcoming data heterogeneity. The participating devices in a federated system are often deployed across a variety of users and environments, resulting in a non-IID data distribution. As the level of heterogeneity intensifies, optimization becomes increasingly difficult. Various techniques have been proposed for alleviating this issue. These primarily consist of modifications to the local or global objectives through proximal terms, regularization, and improved aggregation operations [2, 3, 4, 5, 6]. More recently, some works have investigated the role of model initialization in mitigating such effects [7, 8]. Inspired by the common usage of pre-trained models for facilitating strong transfer learning in centralized training, researchers employed widely available pre-trained weights for initialization in FL and were able to close much of the gap between federated and centralized performance.
Still, while pre-trained initializations are effective for alleviating heterogeneity effects in FL, another key challenge is left unaddressed; that is, communication constraints. This is often the primary bottleneck for real-world federated systems [9]. In the standard FL framework [10], updates for all model parameters are sent back and forth between the server and participating clients each round. This can quickly put a massive communication burden on the system, especially if more capable models beyond very small MLPs are used.
When employing strong pre-trained models, the number of parameters can be large, such as for current state-of-the-art transformers. For example, ViT-Base (ViT-B) [11] has 84 million parameters, let alone the current significant progress in large foundation models (e.g., GPT-4 [12] has 1 trillion parameters). Those large models would simply exacerbate the communication overhead to insurmountable levels. As a compromise, most existing FL work focuses on the performance of smaller Convolutional Neural Networks (e.g., ResNet [13]) on smaller datasets (e.g., CIFAR-10 [14], EMINIST [15]). Considering the thriving progress in large pre-trained Foundation Models [16], an efficient framework enabling these large pre-trained models will be significant for the FL community.
Based on the previous study on centralized training [17, 18, 19, 20], we note that pre-trained models have strong representations, and updating all the weights during fine-tuning is often not necessary. Various parameter-efficient fine-tuning methods (e.g., fine-tuning only a subset of the parameters or the bias terms) for centralized training have been proposed in the literature and show that successful and efficient adaptation is possible, even under domain shift [18, 17, 19]. We reason that such insight is applicable to FL, where each client can be thought of as a shifted domain on which we are fine-tuning. By leveraging pre-trained weights, it may be possible to simply update a small portion of the weights for each client. This will significantly reduce the communication burden on the system, as the updates communicated with the server will consist of just a fraction of the total model parameters.
Can we reap these potential communication benefits while still achieving strong performance in FL? Unfortunately, operating conditions in FL are difficult, requiring successful convergence under varying data heterogeneity levels, random client availability, and differential privacy procedures. Therefore, we are unable to properly assess this possibility of benefit based on existing literature, as the effectiveness of parameter-efficient fine-tuning methods has not been explored in such situations in FL. To fill this gap, we explore the viability of a Federated Parameter-Efficient Fine-Tuning (FedPEFT) framework with a systemic empirical study on a comprehensive set of FL scenarios including communication analysis about communication cost for each method to enable pre-trained models, capability analysis of each method with unlimited communication budget, and robustness analysis of each method when additional constraints (i.e., differential privacy or data scarcity) applied. The framework is illustrated in Fig. 1. We deploy parameter-efficient fine-tuning methods to adapt pre-trained models and enable massive reductions in communication overheads.
The contribution of this paper is summarized as follows:
  • We introduce FedPEFT, a new federated learning framework that simultaneously addresses data heterogeneity and communication challenges. FedPEFT is the first federated learning framework that enables the leveraging of strong pre-trained models in FL while maintaining an extremely low communication cost.
  • We present a systematic study of the FedPEFT framework with various fine-tuning methods under heterogeneous data distributions, client availability ratios, and increasing degrees of domain gap relative to the pre-trained representations on both image and video domains, showing the capability of FedPEFT. (Sections IV-B and IV-C)
  • To ensure FedPEFT is practical for the complex environments of FL, we further analyze the robustness of FedPEFT among low-data regimes, differential privacy operations, and an ablation study with different pre-trained backbones. (Sections IV-D and V-A)
II Related Work
Federated Learning. FL is a decentralized training paradigm composed of two procedures: local training and global aggregation. Therefore, most existing work focuses on either local training [4, 21, 2] or global aggregation [22, 23] to learn a better global model. Another line of work cuts into this problem by applying different initializations to help both procedures. [8] shows that initializing the model with pre-trained weights can make the global aggregation of FedAvg more stable, even when pre-trained with synthetic data. Furthermore, [7] presents the effectiveness of pre-training with different local and global operations. However, these works focus purely on the effect of initialization in a standard FedAvg framework and do not consider the communication constraints of the system. Our work pushes the envelope further by leveraging strong pre-trained models (even large, capable transformers) in FL while effectively handling the communication issue via parameter-efficient fine-tuning.
Communication in Federated Learning. Communication constraints are a primary bottleneck in federated learning. To reduce the communication cost, several previous work leverage model compression techniques [24, 25]. Such works do not change the training paradigm but rather post-process the local model to reduce communication costs. For instance, [24] proposes approaches that parameterize the model with fewer variables and compress the model in an encoding-decoding fashion. However, the minimal requirement to maintain all the information is still high when facing today’s large models. Meanwhile, another line of work changes the training paradigm by learning federated ensembles based on several pre-trained base models [26]. In this way, only the mixing weights of the base models will be communicated in each round. This approach aims to reduce the burden of downloading and uploading the entire model in each round. However, the base models are not directly trained, and the final performance is highly related to the base models. Meanwhile, model ensembles will take more time and space, which is often limited on the client side. Our framework follows the strategy of this line of work that does not transmit the entire model, but we use only one pre-trained model instead of several base models and only transmit a subset of the parameters instead of the model ensembles. Therefore, no additional time or space is required.
Parameter-Efficient Fine-tuning. Fine-tuning is a prevalent topic in centralized transfer learning, especially in this era of the “Foundation Model” [16]. A significant line of work is to reduce the trainable parameter number, i.e., parameter-efficient fine-tuning (PEFT) [20, 27, 28, 29, 30, 31, 17, 32]. PEFT has emerged as a pivotal area of research in the field of natural language processing (NLP) [19, 33, 34, 35] and further adapted into more fields such as computer vision (CV) [18, 36, 37, 38, 39, 40, 41]. With different focuses, PEFT methods can be divided into three categories: 1) Input Adaptation methods such as prompt-tuning [18, 30] focusing on adding learnable context to the input data, 2) Backbone Adaptation methods such as adapter-tuning [38, 19, 33], and 3) Specification methods such as bias-tuning [17, 31]. PEFT enables easier access and usage of pre-trained models by reducing the memory cost needed to conduct fine-tuning due to fewer computed gradients. In federated learning, PEFT has an additional benefit that is not salient in centralized training: reducing communication costs. By introducing PEFT to federated learning, our work can take advantage of a strong (and even large) pre-trained model while significantly reducing communication costs. Several concurrent and follow-up works study the PEFT methods under FL settings in NLP [42, 43, 44, 45]. Complementarily, our framework is the first work to present a comprehensive study of PEFT in various FL settings in computer vision under both image and video tasks.
III Federated Parameter-Efficient Fine-Tuning
III-A Problem Formulation
In this section, we formally describe the federated learning objective and federated parameter-efficient fine-tuning. Using a classification task as an example, $K$ samples in a dataset ${\mathbb{D}}=\{({\bm{x}}_{k},y_{k})_{k=1}^{K}\}=\cup_{n=1}^{N}{\mathbb{D}}_{n}$, where ${\bm{x}}$ is the input and $y\in\{0,1,\dots,C-1\}$ is the label, are distributed among $N$ clients. Each client has a local model $\{\phi_{i}\}_{i=1}^{N}$ parameterized by $\{{\bm{\theta}}_{i}\cup\delta_{i}\}_{i=1}^{N}$, where ${\bm{\theta}}$ is the pre-trained weights and $\delta$ is the trainable parameters. The goal of federated learning is to learn a global model $\phi$ parameterized by ${\bm{\theta}}\cup\delta$ on the server from $M$ sampled client models in $T$ communication rounds by minimizing the global objective $F$ as
$\min_{\delta}F(\phi)=\frac{1}{|{\mathbb{D}}_{test}|}\sum_{i=1}^{|{\mathbb{D}}_{test}|}\ell(y_{k},\phi_{i}^{(t)}({\bm{x}}_{k}))$ (1)
on a hold-off test set ${\mathbb{D}}_{test}$ with a loss function $\ell$. Compared with traditional FL updating the entire $\phi$, only $\delta$ is updated in FedPEFT.
At the beginning of training, ${\bm{\theta}}^{(0)}$in the global model $\phi^{(0)}$ is initialized with pre-trained weights, and $\delta^{(0)}$ is randomly initialized, where the superscript $t$ indicates the model at round $t$. In each round $t$, $M$ clients will be selected for communication, and $\{\phi_{i}^{(t)}\}_{i=1}^{M}$ will be initialized by $\phi^{(t)}$ and updated by
$\min_{\delta_{i}}F_{i}(\phi_{i}^{(t)})=\frac{1}{|{\mathbb{D}}_{i}|}\sum_{k=1}^{|{\mathbb{D}}_{i}|}\ell(y_{k},\phi_{i}^{(t)}({\bm{x}}_{k}))$ (2)
for $E$ epochs, where $F_{i}$ is the local objective. After the local updates, the server will receive and aggregate the trainable parameters $\{\delta_{i}^{(t)}\}_{i=1}^{M}$ with the FedAvg algorithm to a new global model
$\phi^{(t+1)}=\sum_{m=1}^{M}\frac{|{\mathbb{D}}_{m}|}{\sum_{i=1}^{M}|{\mathbb{D}}_{i}|}\phi_{i}^{(t)}.$ (3)
This procedure is repeated from $t=0$ to $t=T-1$. During the client-server communication, we only take the communication cost for the model into consideration, assuming the remaining communication costs are fixed. Therefore, the communication cost $C$ is proportional to the transmission parameters number, thus can be formulated as
$C\propto|\delta|\cdot M,$ (4)
We take the one-way communication cost (i.e., upload or download) as the metric. The final goal of this problem is to minimize the $C$ while maintaining server accuracy.
Algorithm 1 Algorithm of FedPEFT framework
Input: $N$ clients indexed by $i$, participating-client number $M$, communication rounds $T$, trainable parameters $\delta$ of the model where $|\delta|<<|\phi|$, pre-trained model weights ${\bm{\theta}}$, random initialized $\delta^{(0)}$, and local epoch number $E$. Server executes:       initialize $\phi^{(0)}$ and $\{\phi_{i}^{(0)}\}_{i=1}^{N}$ with ${\bm{\theta}}$ and $\delta^{(0)}$       for each round $t=0,2,...,T-1$ do        ${\mathbb{S}}_{t}\leftarrow$ (random set of $M$ clients)        for each client $i\in{\mathbb{S}}_{t}$ in parallel do         $\delta_{i}^{(t)}\leftarrow\delta^{(t)}$         $\delta_{i}^{(t+1)}\leftarrow\textbf{ClientUpdate}(\delta_{i}^{(t)},i)$        $\delta^{(t+1)}=\sum_{i=1}^{M}\frac{|{\mathbb{D}}_{i}|}{\sum_{i=1}^{M}|{\mathbb{D}}_{i}|}\delta_{i}^{(t+1)}$        $\phi^{(t+1)}\leftarrow\{{\bm{\theta}},\delta^{(t+1)}\}$       return $\phi^{(T)}$ ClientUpdate ($\delta,i$):       $\delta\leftarrow$ perform local training on $\delta$ with ${\mathbb{D}}_{i}$ for $E$ epochs       return $\delta$
Figure 2: Methods to fine-tune each layer in a pre-trained backbone, where $h$ means the input, $\phi$ means the pre-trained layer, and $\phi_{w},\phi_{b}$ mean its weight and bias parameters, respectively. (a) Fully fine-tune the entire model. (b) Only tune the classification head with the frozen backbone. Blue and red rectangles mean frozen and trainable parameters, respectively. (c) Only tune the bias term of each layer. (d) Concatenate learnable parameters $\phi_{p}$ as prompts to the input in each layer and only tune the prompts with the frozen backbone. (e) Add an adapter $\phi_{a}$ to each layer of the backbone, and only tune the adapters with the frozen backbone.
III-B FedPEFT
In conventional federated learning, updates for the entire model need to be repeatedly sent to and from the server, resulting in significant communication costs, especially when larger, more capable modern neural network architectures are employed. To reduce this heavy burden, we deploy parameter-efficient fine-tuning methods to adapt pre-trained models to the local clients rather than fully fine-tuning all parameters, which is described in Algorithm 1. In the FedPEFT framework, illustrated in Fig. 1, only a small amount of parameters in the local model will be downloaded, trained, and uploaded in each communication round. For instance, FedPEFT reduces the size of communication each round from 328MB (85.88M parameters)/Client to 0.68MB (0.17M parameters)/Client when using a pre-trained ViT-Base as the backbone in our default setting introduced in Section IV-A.
To implement FedPEFT, we provide a canonical baseline approach (Head-tuning) and three prototypes leveraging different parameter-efficient fine-tuning methods (Bias, Adapter, and Prompt), which are detailed in the following.
To reduce the number of trainable parameters, one intuitive method, Head-tuning, is to freeze the backbone $\phi$ and only train the head $c$. This method is historically the most common fine-tuning procedure, and therefore we use it as a baseline for FedPEFT. However, the adaptation ability of this method is limited, as no adjustment is made to the network representation prior to the final output head. This can be problematic in the presence of a more intense domain shift. Therefore, we consider the following approaches as primary prototypes for FedPEFT:
FedPEFT-Bias. Bias-tuning [17] aims to adapt the pre-trained backbone by only fine-tuning a specific group of parameters, the bias term. In this way, the backbone can be trained with moderate adjustments to prevent damaging the upstream representation. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
[…]
where $\phi_{w}^{L}$ and $\phi_{b}^{L}$ are weight and bias of $\phi_{L}$, blue parameters are frozen, while red parameters are trainable.
FedPEFT-Adapter. Instead of directly tuning existing parameters in the backbone like Bias-tuning, Adapter-tuning [19] adds a few parameters called adapters inside the backbone $\phi$ instead. Usually, adapters will be deployed in each layer of the backbone to perform transformations on different levels of the pre-trained feature while the backbone stays frozen. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
[…]
where $\phi_{a}$ is the adapter.
FedPEFT-Prompt. Prompt-tuning [18] takes a slightly different approach from the other fine-tuning methods. Specifically, it concatenates trainable parameters, called prompt embeddings, to the input embedding and hidden states in each layer. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
[…]
We illustrate the differences between all baseline and prototype methods in Fig. 2. Besides the above prototypes, our framework is compatible with other PEFT methods. We provide additional discussion on Sec. V-B.
III-C Convergence Guarantee
Based on the convergence of FedAvg in [1, 3, 46], in this section, we will comment on the convergence of FedPEFT. For ease of notation, we consider, at each round, $S$ to be the number of clients sampled. We require the following assumptions.
Assumption 1
(Global minimum) For the global objective $F$, there exists $\phi^{\star}$ such that, $F(\phi^{\star})=F^{\star}\leq F(\phi)$, for all $\phi\in\mathbb{R}^{d}.$
Assumption 2
($\beta$-Smoothness) The loss function $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ at each node is $\beta$-smooth, i.e. $f_{i}(y)\leq f_{i}(x)+\nabla f_{i}(x)^{\top}(y-x)+\frac{\beta}{2}\|y-x\|^{2}$ for all $x,y\in\mathbb{R}^{d}$.
Remark 1
(PEFT-FT gap) The above assumption implies that $F$ is $\beta$-smooth. Therefore, the gap between PEFT and Full Fine-tuning can be proved as $|F({\bm{\theta}}^{(T)}\cup\delta^{(0)})-F({\bm{\theta}}^{(T)}\cup\delta^{(0)})={\mathcal{O}}(\frac{\beta}{2}(\|{\bm{\theta}}^{(T)}-{\bm{\theta}}^{(0)}\|^{2}+\|\delta^{(T)}-\delta^{(0)}\|^{2}))|$.
Assumption 3
There exist constants $G\geq 0,B\geq 1$, such that for all $x\in\mathbb{R}^{d}$, the stochastic noise, $\xi_{i,t}$ follows
$\frac{1}{N}\sum_{i=1}^{N}\|\nabla f_{i}(x)\|^{2}\leq G^{2}+B^{2}\|F(x)\|^{2}.$
Assumption 4
(Bounded variance) Let $g_{i}(\phi):=\nabla f_{i}(\phi,z_{i(k)})$ be the unbiased stochastic gradient of $f_{i}$ with bounded variance. That is, there exists, $\sigma\geq 0$ such that, $\mathbb{E}_{z_{i(k)}}\left[\|g_{i}(\phi)-\nabla f_{i}(\phi)\|^{2}\right]\leq\sigma^{2},$ for all $\phi,i$, where $z_{i(k)}$ is the $k^{\rm th}$ sample data at the $i^{\rm th}$ client.
Based on the vanilla FedAvg framework, we can give our main convergence result as
Theorem 1
Let $F$ satisfies Assumptions 1-4. Then
$\mathbb{E}\left[\|\nabla F(\phi^{(T)})\|^{2}\right]\leq{\mathcal{O}}\left(\frac{\beta\sqrt{(F(\phi^{(0)})-F^{\star})}}{\sqrt{TEM}}+P\right),$
where $P=\frac{\beta}{2}(\|{\bm{\theta}}^{(T)}-{\bm{\theta}}^{(0)}\|^{2}+\|\delta^{(T)}-\delta^{(0)}\|^{2})$.
III-D Privacy Discussion
Federated learning inherently ensures data privacy, as it keeps the training data localized. Consequently, our proposed FedPEFT framework, by design, does not introduce any additional risk of privacy leakage beyond what is intrinsic to FL itself. However, it is crucial to acknowledge the vulnerability of FL systems to gradient inversion attacks [47, 48], where an adversary could potentially reconstruct original training data from shared gradients. This type of attack typically requires smaller batch sizes to be effective, as the precision of the information contained within the gradients diminishes with larger batch sizes, significantly reducing the feasibility of such attacks [49]. In light of this, FedPEFT inherently encourages the use of larger batch sizes compared to conventional full fine-tuning methods, as FedPEFT requires gradients for much fewer parameters and, therefore, very little memory cost during training. Enabling larger batch sizes not only optimizes the training efficiency but also fortifies the privacy-preserving nature of the federated learning framework, further mitigating the risks associated with gradient inversion attacks.
TABLE I: Communication analysis on the image domain. The communication cost is computed with 4B/parameter. The averaged final accuracy, i.e., $t=T=50$, and the standard deviation of three different random seeds are reported for each data set. The number of tuned parameters is computed based on CIFAR-100, but it may be slightly different for each dataset. The first section shows the change in accuracy when decreasing the participating client number. The gray numbers indicate the baseline performance with no decrease in the participating client number. The second section shows the change in accuracy when we reward the low communication cost of head-tuning by increasing the number of participating clients. The third section shows the accuracy when we fully fine-tune a lightweight model, ShuffleNet V2 0.5×0.5$\times$0.5 × [50]. The fourth section shows the performance of each prototype of FedPEFT.
Model Method # Tuned Params $\times$ Clients Comm. Cost Resisc45 CIFAR-100 PCam
ViT-B Full Fine-tuning 85.88M $\times$ 8 2.56GB 91.49$\pm$0.82 91.73$\pm$0.43 85.41$\pm$2.41
ViT-B Full Fine-tuning 85.88M $\times$ 4 1.28GB 92.13$\pm$0.87 89.69$\pm$0.30 81.93$\pm$3.54
ViT-B Full Fine-tuning 85.88M $\times$ 2 656MB 87.68$\pm$1.32 87.03$\pm$0.18 82.20$\pm$1.22
ViT-B Full Fine-tuning 85.88M $\times$ 1 328MB 73.38$\pm$1.95 74.79$\pm$0.77 80.18$\pm$1.83
ViT-B Head-tuning 0.08M $\times$ 8 2.44MB 77.30$\pm$1.03 72.45$\pm$0.08 74.82$\pm$2.40
ViT-B Head-tuning 0.08M $\times$ 64 19.53MB 83.58$\pm$0.45 75.45$\pm$0.16 77.82$\pm$0.37
ShuffleNet Full Fine-tuning 0.44M $\times$ 8 13.43MB 63.52$\pm$0.50 49.81$\pm$1.94 76.52$\pm$3.35
ViT-B FedPEFT-Bias 0.18M $\times$ 8 5.49MB 89.04$\pm$0.80 90.79$\pm$0.25 85.51$\pm$0.66
ViT-B FedPEFT-Adapter 0.23M×$\times$ 8 7.02MB 87.20$\pm$0.78 87.74$\pm$0.55 78.67$\pm$1.85
ViT-B FedPEFT-Prompt 0.17M×$\times$ 8 5.19MB 83.35$\pm$0.76 89.78$\pm$0.84 86.50$\pm$0.85
TABLE II: Communication analysis on the video domain.
Model Method # Tuned Params $\times$ Clients Comm. Cost UCF101 HMDB51 UCF-CRIME
ViT-B Full Fine-tuning 86.30M $\times$ 4 1.29GB 94.22$\pm$0.23 70.34$\pm$0.28 34.37$\pm$0.76
ViT-B Full Fine-tuning 86.30M $\times$ 2 656MB 93.85$\pm$0.33 69.06$\pm$0.48 32.03$\pm$0.26
ViT-B Full Fine-tuning 86.30M $\times$ 1 328MB 92.61$\pm$0.16 59.88$\pm$0.27 24.21$\pm$0.26
ViT-B Head-tuning 0.08M $\times$ 4 1.22MB 88.57$\pm$0.30 64.98$\pm$0.32 32.03$\pm$0.76
ViT-B Head-tuning 0.08M $\times$ 32 9.76MB 89.74$\pm$0.45 63.87$\pm$0.66 32.81$\pm$0.76
X3D-S Full Fine-tuning 3.07M $\times$ 4 47.96MB 36.68$\pm$1.67 27.74$\pm$0.24 21.42$\pm$0.76
ViT-B FedPEFT-Bias 0.18M $\times$ 4 2.75MB 92.34$\pm$0.30 69.37$\pm$0.32 34.38$\pm$0.52
ViT-B FedPEFT-Adapter 0.23M×$\times$ 4 3.51MB 92.82$\pm$0.44 69.96$\pm$0.24 33.76$\pm$0.26
ViT-B FedPEFT-Prompt 0.17M×$\times$ 4 2.60MB 93.82$\pm$0.66 70.87$\pm$0.16 34.38$\pm$0.26
IV Experiments
The capability of full fine-tuning in terms of accuracy has been illustrated in recent work [7, 8]. Thus we regard it as a competitive baseline for FedPEFT. To verify the performance of FedPEFT comprehensively, we evaluate the server accuracy with each method from three perspectives and aim to answer the following questions:
Communication Analysis: When faced with a limited communication budget, there are several conventional solutions to reduce costs, e.g., sampling fewer clients each round or using a lightweight model. Can FedPEFT outperform other solutions in terms of communication cost and accuracy? (RQ1)
Capability Analysis: When the communication budget is amply sufficient for all approaches, we want to evaluate the trade-off of training fewer parameters with FedPEFT. Can FedPEFT outperform full fine-tuning and training from scratch within various federated learning settings and increasing levels of downstream domain gap? (RQ2)
Robustness Analysis: In a lot of application scenarios, there will be additional challenges for FL, such as privacy-preserving requirements (i.e., differential privacy) and data scarcity (i.e., very small amount of data on each client). We want to evaluate the robustness of each method under such scenarios. Can FedPEFT outperform full fine-tuning in terms of robustness? (RQ3)
IV-A Experiments details
Dataset. For our study, we focus on computer vision (CV) applications as our testbed. Specifically, we investigate the performance of each method on the Image and Video domains with image classification and action recognization tasks. For the image classification, we employ ImageNet-21K [51] as the pre-training dataset, which is a primary dataset for pre-training models in CV. Then we select three datasets for the downstream tasks that have increasing degrees of domain gap compared to ImageNet-21k, and we visualize and quantify the domain gap in Section IV-C: Resisc45 [52], CIFAR-100 [14] and PCam [53]. Resisc45 is a remote sensing image dataset containing 18,900 training samples and 6,300 testing samples for 45 scene classes, which has a small domain gap with ImangeNet-21K. CIFAR-100 contains 50,000 training samples and 10,000 testing samples for 100 classes of nature objects (e.g., sunflowers, buses), which has a medium domain gap. PCam is a medical dataset containing 32,768 colored images extracted from histopathologic scans of lymph node sections, which has a much larger domain gap. Each image is annotated with a binary label indicating the presence of metastatic tissue. We sample 20,000 images from the original train set for training and use the entire test set for evaluation. We use the official splitting for CIFAR-100 and PCam and a public splitting from AiTLAS [54] for Resisc45.
For action recognition evaluation, we employ Kinetics-400 [55] as the pre-training dataset, which is a common dataset for pretraining in action recognition. We select three datasets with varying degrees of domain gap as compared to Kinetics-400: UCF101 [56], HMDB51 [57] and UCF-CRIME [58]. The UCF101 [56] dataset contains 13,320 videos from 101 classes. The HMDB51 [57] dataset contains 6849 videos divided into 51 categories based on everyday human actions. UCF-CRIME [58] consists of surveillance videos with 13 anomalous activities as action recognition classes. For UCF-CRIME, we use a train/val split of 40/10 for each class.
Figure 3: Server accuracy given the total communication budget. The communication cost is computed with 4B/parameter, and the max number of communication rounds is 50. The number in the bracket next to the method indicates the number of participating clients $m$. The transparency of the line indicates the ratio between $m$ and total client number $N=64$. The horizontal dashed line shows a target accuracy of $85\%$.
Experimental Setting. Our default experimental setting is to split the dataset across $N=64$ clients and sample $M=8$ clients each round. The global aggregation will be performed after $E=10$ local epochs. A total of $T=50$ rounds of communication will be performed. To simulate heterogeneous data, we partition samples in each class to all clients following a Dirichlet distribution, as common in the literature [4, 5, 21], with $\alpha=0.1$ for CIFAR-100 and Resisc45 and $\alpha=0.5$ for PCam based on the class number. Any modifications to this setting in subsequent experiments will be clearly noted.
For each action recognition experiment, the data is split across $N=32$ clients and sample $M=4$ clients each round with constant $\alpha=0.1$ across all three video datasets.
Implementation Detail. We choose ViT-B [11] with image size 224 and patch size 16 as our backbone. For the image domain, the backbone is pre-trained on ImageNet-21K [51], as available in the timm library [59]. The images for the downstream datasets are resized to $224\times 224$. Images from CIFAR-100 are augmented by random cropping with a padding of $4$ and random horizontal flipping, and Resisc45 and PCam are augmented only with random horizontal flipping. For the video domain, we choose vanilla ViT-Base (ViT-B/16) [11] with joint space-time attention as our backbone model [60] using VideoMAE [61] pre-trained weights on Kinetics-400 [55]. We perform the experiments on 8 Nvidia RTX A5000 GPUs with a batch size of $64$. All reported numbers are run multiple times and averaged. The base hyperparameters for each method are described below.
  • Full fine-tuning: We use SGD [62] optimizer with learning rate $0.001$ and weight decay $0.0001$.
  • Head-tuning: We use one linear layer as the classification head and use the SGD optimizer with learning rate $0.005$ and weight decay $0.0001$.
  • FedPEFT-Bias: There is no additional hyperparameter for Bias. We use SGD optimizer with learning rate $0.01$ and weight decay $0.0001$.
  • FedPEFT-Adapter: We use a Bottleneck Adapter with residual connections as the adapter. We insert the adapter to each layer of the backbone after the feed-forward block following [19] with a reduction factor of $8$ and the GELU as the activation function. We use SGD optimizer with learning rate $0.005$ and weight decay $0.0001$.
  • FedPEFT-Prompt: We follow the design of VPT-Deep [18], and use a prompt length as $10$ to prepend prompt tokens to the input and the hidden embedding in each layer. We use SGD optimizer with learning rate $0.01$ and weight decay $0.0001$.
IV-B RQ1: Communication Analysis
To verify the effectiveness of FedPEFT and answer the first research question (RQ1, Section IV), we compare it with three baselines while monitoring the communication budget: a) Full fine-tuning of our default model (ViT-B). We vary the number of participating clients to show different levels of communication requirements. b) Head-tuning. The communication cost of head-tuning is naturally lower than other methods, so we increase the participating clients to make it a stronger baseline. c) Fully fine-tune a light-weighted model (ShuffleNet V2 0.5$\times$ [50] for images and X3D-S [63] for videos) with a similar parameter scale.
As demonstrated in TABLE I for the image domain and TABLE II for the video domain, all FedPEFT methods achieve better results in many cases compared with other approaches, even with significantly fewer communicated parameters. We find that full fine-tuning needs several orders of magnitude of communication to achieve a comparable result with FedPEFT. For instance, it needs at least 187$\times$ and 477$\times$ more parameters to reach and outperform FedPEFT on CIFAR-100. Interestingly, full fine-tuning performs well on Resisc45 where the domain gap is smaller, even when the participating-client number is low. However, when the domain gap increases, more participating clients will be needed to outperform FedPEFT, and finally it fails to outperform FedPEFT even without reducing the participating-client number on PCam where a large domain gap exists. Meanwhile, head-tuning lags behind most other approaches, but the performance is stable with different levels of domain gap, while the ShuffleNet model only achieves $71\%$, $55\%$, and $89\%$ of accuracy on Resisc45, CIFAR-100, and PCam with 2.4$\times$ the communication cost compared with FedPEFT-Bias. Besides, the standard deviation of FedPEFT is lower than most other solutions, especially when the domain gap is large, showing the stability of FedPEFT. For the video domain, the conclusion is consistent.
In Fig. 3, we also report the server accuracy that can be achieved for each method given the communication budgets using CIFAR-100 as an example. The communication cost per communication round for full fine-tuning is even higher than the total communication cost for FedPEFT to converge to similar final server accuracy. Meanwhile, all FedPEFT prototypes only require megabytes level communication, while full fine-tuning requires gigabytes level communication to reach a given target accuracy (e.g., 85% in Fig. 3), showing the efficiency of FedPEFT. For the inter-prototype comparison, FedPEFT-Bias stands out for its highest efficiency. We provide further discussions on the performance of each prototype in Section IV-C.
(a) Extracted feature. The number in the bracket is the domain gap to the pre-training dataset.
(b) Relative accuracy of each method compared with full fine-tuning with different domain gaps.
Figure 4: Visualization and analysis of domain gap. (a) We apply Linear Discriminant Analysis (LDA) to the extracted feature from the pre-trained backbone to visualize and quantify the domain gap. (b) We plot the performance (relative accuracy) of each PEFT method compared against full fine-tuning. The gray dashed line indicates the performance of full fine-tuning (i.e., reference), and the three points in each line are the relative server accuracy on Resisc45, CIFAR-100, and PCam, respectively.
TABLE III: Capability analysis among different federated learning settings on CIFAR-100 and UCF-101. Training from scratch in more complicated settings will lead to a lower result than in the first setting, which is omitted here. Bold-style shows the best performance among all methods or among prototypes in FedFEFT.
Client Method Image Video
# Tuned Params Homogeneous Heterogeneous # Tuned Params Homogeneous Heterogeneous
[…]
Full Fine-tuning 85.88M 93.70 93.50 86.30M 95.32 95.17
Head-tuning 0.08M 78.11 77.59 0.08M 89.90 85.44
FedPEFT-Bias 0.18M 91.89 90.25 0.18M 93.14 92.88
FedPEFT-Adapter 0.23M 90.21 88.77 0.23M 93.73 93.48
FedPEFT-Prompt 0.17M 92.09 90.37 0.17M 94.23 94.09
[…]
Head-tuning 0.08M 76.65 62.80 0.08M 89.68 84.32
FedPEFT-Bias 0.18M 91.35 86.18 0.18M 93.05 92.64
FedPEFT-Adapter 0.23M 89.48 80.08 0.23M 93.53 93.27
FedPEFT-Prompt 0.17M 91.60 85.54 0.17M 94.11 93.89
[…]
Head-tuning 0.08M 78.45 75.51 0.08M 59.56 52.74
FedPEFT-Bias 0.18M 92.71 91.71 0.18M 78.21 75.33
FedPEFT-Adapter 0.23M 90.50 89.26 0.23M 78.41 75.88
FedPEFT-Prompt 0.17M 91.87 90.96 0.17M 78.69 76.11
[…]
Head-tuning 0.08M 77.59 72.55 0.08M 56.54 46.62
FedPEFT-Bias 0.18M 92.49 91.02 0.18M 77.19 74.42
FedPEFT-Adapter 0.23M 90.39 88.05 0.23M 77.69 75.18
FedPEFT-Prompt 0.17M 92.00 89.90 0.17M 78.56 75.38
IV-C RQ2: Capability Analysis
To study and understand our second research question (RQ2, Section IV), we analyze the impact of the domain gap between the model pre-training dataset and the dataset for FL (Section IV-C) and systemically perform experiments on CIFAR-100 across different federated learning scenarios by varying client status and data distribution (Section IV-C).
Capability with Domain Gap. Domain gap is a realistic concern when deploying pre-trained models for downstream tasks. As shown in TABLES I and II, the performance on different datasets varies a lot. Besides the difference in the difficulty between all datasets, the domain gap to the pre-trained dataset is also a key concern here. To further discuss the impact of the domain gap for each method, we use our default setting shown in Section IV-A on the image domain as an instance to visualize and quantify the domain gap between each downstream dataset and the pre-training dataset. Specifically, we adapt Linear Discriminant Analysis (LDA) for all extracted features for the test samples in each dataset from the pre-trained backbone to reduce the dimension. We compute the center of each dataset and then compute the distance to the center of the pre-training dataset as the quantifying result of the domain gap, as shown in Fig. 4a.
In Fig. 4b, we present the performance of all approaches with an increasing degree of domain gap compared to the ImageNet-21k pre-training dataset. Interestingly, full fine-tuning falls further behind as the data domain gap widens in the PCam scenario, largely unable to keep up with FedPEFT despite requiring a massive communication budget. This phenomenon when a pre-trained model meets out-of-domain data has been studied under centralized settings [64]. It was found that the pre-trained upstream representations are still meaningful even with a domain gap. Therefore, fully fine-tuning the backbone with out-of-domain data can damage the high-level semantics inside these upstream representations due to overfitting, especially when the data size is small. This is particularly relevant in FL, where overfitting and subsequent client drift [65, 4, 3] are prone to occur.
On the opposite end of the spectrum, by not fine-tuning the backbone at all, head-tuning maintains similar accuracy despite the domain gap. This shows the robustness of the pre-trained high-level semantics across domains, supporting the conclusion that there is meaningful high-level semantics inside of the upstream representations.
Still, the tight restriction on head-tuning is perhaps a bit too far, as the accuracy on all datasets is still low overall. Between the two extremes of head-tuning and full fine-tuning, FedPEFT approaches may be able to suitably adapt the upstream representations without excessively damaging them. Specifically, FedPEFT-Bias operates with parameter-level control for each parameter pair containing weight and bias terms. The representation can then preserve the high-level semantics by freezing the weight term (maintaining the direction in the feature space) and still adapting via the bias term (shifting in the feature space). FedPEFT-Adapter and FedPEFT-Prompt have slightly different mechanisms, controlling the backbone by transforming the intermediate hidden representations via adapters and prompts. Specifically, FedPEFT-Prompt adds additional hidden states before the original hidden states without changing the original representation, while FedPEFT-Adapter transforms hidden states into a new space. Consequently, FedPEFT-Prompt shows stronger robustness in handling larger domain gaps than FedPEFT-Adapter. Of these approaches, FedPEFT-Prompt is the most stable under the domain gap, surpassing full fine-tuning by  1.1% on PCam. Overall, we hypothesize that more fine-tuning freedom will be better when the domain gap is minor, but moderate fine-tuning is needed to maintain, as well as control, the high-level semantics when the domain gap is large.
TABLE IV: Robustness analysis for privacy-preserving. The red number indicates the performance difference under DP.
Method Image Video
w/o DP w/ DP w/o DP w/ DP
Full Fine-tuning 92.09 77.61 (-14.48) 94.22 86.34 (-7.88)
Head-tuning 72.55 62.20 (-10.35) 88.57 83.09 (-5.48)
FedPEFT-Bias 91.02 84.98 (-6.04) 92.34 86.61 (-5.73)
FedPEFT-Adapter 88.05 79.05 (-9.00) 92.82 85.57 (-7.25
FedPEFT-Prompt 89.90 78.35 (-11.55) 93.82 87.64 (-6.18)
Capability with Different FL Settings. In application scenarios, the setting of federated learning can vary substantially. It is important to show the capability to maintain high performance in diverse settings. We present results for all approaches with different client availability ratios and data distributions in TABLE III and draw the following conclusions from the experiments:
First, we see that fine-tuning the pre-trained model shows a significant improvement over training from scratch, especially in heterogeneous scenarios. This finding is in agreement with other very recent works [8, 7], which note the stabilization effect of pre-trained initialization in federated optimization. When only fine-tuning the head, the performance is still much better than training the entire model from scratch (Scratch) but remains low in comparison to other methods across all settings. We again find that head-tuning simply lacks adaptation ability, holding too closely to the upstream representation.
On the other hand, we find that FedPEFT achieves comparable results ($\mathbf{\geq 95\%}$) to full fine-tuning with less than 0.3% of the trainable parameters. This ability to maintain accuracy performance in various scenarios is crucial for FL, as oftentimes, the exact setting and distributions are not known ahead of time. Meanwhile, for inter-prototype comparison, we find that FedPEFT-Bias outperforms other prototypes in almost all settings in the image domain, while FedPEFT-Prompt shows leading performance among all prototypes in the video domain. This provides guidance in application to choose the prototype.
IV-D RQ3: Robustness Analysis
In this section, we further investigate our third research question (RQ3) in two critical FL scenarios evaluated by CIFAR-100 and UCF-101.
Differential Privacy. A fundamental property of federated learning is privacy protection. However, various works [47, 48] have demonstrated how the client data can be reconstructed from the raw gradient updates received by the server in some scenarios. To protect client data privacy from such attacks, differential privacy (DP) [9, 66, 67, 68] has become standard practice. Therefore, we first study FedPEFT and other baselines under DP.
TABLE V: Robustness analysis for data scarcity. $K$ indicates the total sample number of all clients.
Method $K=1000$ $K=1500$ $K=2000$
Image Video Image Video Image Video
Full Fine-tuning 66.52 87.50 67.47 88.34 77.67 90.54
Head-tuning 52.13 83.46 56.52 85.56 60.15 86.76
FedPEFT-Bias 76.40 85.34 81.14 87.36 83.83 88.85
FedPEFT-Adapter 71.34 86.17 76.91 88.12 79.22 90.28
FedPEFT-Prompt 63.77 87.46 71.94 88.22 76.89 90.45
To integrate DP, we apply a Gaussian mechanism within the local optimization of each iteration [68] with $\varepsilon=5$ and $\delta=0.001$. We maintain the remaining FL settings as described in Section IV-A, and show the results in TABLE IV. Interestingly, when comparing all methods, full fine-tuning experiences the sharpest drop with DP. This causes its accuracy to fall lower than all the FedPEFT prototypes. To understand this effect, we note that DP applies noise to all trainable parameter gradients. Full fine-tuning, therefore, requires such noise on all model parameters, resulting in a more pronounced negative effect on final performance. On the other hand, the other fine-tuning methods maintain some part of the backbone frozen and have significantly fewer trainable parameters on which adding noise is necessary, limiting the performance drop. Overall, FedPEFT allows for stronger accuracy in DP-enabled federated systems than even full fine-tuning while still maintaining extremely low communication needs.
Data Scarcity. We explore another common yet challenging robustness condition in FL; that is when very little data is available on individual clients. Such data scarcity scenarios are even a tricky problem in centralized training. For example, [69] show that fewer training data will incur damage to the pre-trained representation due to overfitting. In our evaluation for FL, we reduce the total sample number $K$ to $1000$, $1500$, and $2000$. As shown in TABLE V, we find that FedPEFT outperforms full fine-tuning and head-tuning under such low-data scenarios, further revealing its capability to appropriately adapt pre-trained representations to the FL task at hand.
For the inter-prototype comparison, FedPEFT-Bias and FedPEFT-Prompt remain leading the performance in image and video domains, consistent with the conclusion in common scenarios, showing their robustness.
TABLE VI: Ablation study on network backbone and pre-training dataset. 1K and 21K mean pre-trained on ImageNet1K and ImageNet21K respectively. The server accuracy (acc) on CIFAR-100 is reported.
Method Backbone Pre-training Server Acc (%)
Full Fine-tuning ViT-B Supervised (21K) 92.09
ViT-B Supervised (1K) 92.09
ViT-S Supervised (21K) 89.33
ViT-B DINO (1K) 85.62
Head-tuning ViT-B Supervised (21K) 72.55
ViT-B Supervised (1K) 80.78
ViT-S Supervised (21K) 74.40
ViT-B DINO (1K) 77.17
FedPEFT-Bias ViT-B ImageNet-21K 91.02
ViT-B Supervised (1K) 90.73
ViT-S Supervised (21K) 88.40
ViT-B DINO (1K) 85.34
FedPEFT-Adapter ViT-B ImageNet-21K 87.99
ViT-B Supervised (1K) 88.69
ViT-S Supervised (21K) 84.15
ViT-B DINO (1K) 83.29
FedPEFT-Prompt ViT-B ImageNet-21K 89.63
ViT-B Supervised (1K) 89.29
ViT-S Supervised (21K) 87.39
ViT-B DINO (1K) 84.25
TABLE VII: Ablation study on network backbone and pre-training dataset. The server accuracy on UCF-101 is reported.
Method Backbone Pre-training Server Acc (%)
Full Fine-tuning ViT-B Supervised (K400) 94.22
ViT-B Supervised (SSV2) 93.94
ViT-S Supervised (K400) 93.03
Head-tuning ViT-B Supervised (K400) 88.57
ViT-B Supervised (SSV2) 88.12
ViT-S Supervised (K400) 87.64
FedPEFT-Bias ViT-B Supervised (K400) 92.34
ViT-B Supervised (SSV2) 92.16
ViT-S Supervised (K400) 92.10
FedPEFT-Adapter ViT-B Supervised (K400) 92.82
ViT-B Supervised (SSV2) 91.97
ViT-S Supervised (K400) 91.48
FedPEFT-Prompt ViT-B Supervised (K400) 93.82
ViT-B Supervised (SSV2) 93.22
ViT-S Supervised (K400) 91.66
TABLE VIII: Ablation study on different federated learning algorithms. The server accuracy on CIFAR-100 and UCF-101 is reported for the image and video domain, respectively.
Method Image Video
FedAVG FedProx MOON FedAVG FedProx MOON
Full Fine-tuning 92.09 92.11 91.99 94.22 94.44 94.30
Head-tuning 72.55 72.55 72.55 88.57 88.57 88.57
FedPEFT-Bias 91.02 91.02 91.15 92.34 92.30 92.34
FedPEFT-Adapter 88.05 88.19 88.13 92.82 92.96 93.08
FedPEFT-Prompt 89.90 90.20 90.25 93.82 93.92 93.98
TABLE IX: Compatibility of FedPEFT with other PEFT methods. The server accuracy on CIFAR-100 and UCF-101 is reported for the image and video domain, respectively.
Method # Trainable Params Image Video
Full Fine-tuning 85.88M 92.09 94.22
Head-tuning 0.08M 72.55 88.57
FedPEFT-Bias 0.18M 91.02 92.34
FedPEFT-Adapter 0.23M 88.05 92.82
FedPEFT-Prompt 0.17M 89.90 93.82
FedPEFT-Prefix [34] 0.08M 86.26 87.68
FedPEFT-LoRA [29] 0.22M 91.06 92.54
IV-E Insights and Takeaways
Our research findings contribute significant insights to leverage parameter-efficient fine-tuning in federated learning, aiming at reducing communication costs.
  • FedPEFT stands out under stringent communication budgets by offering significant advantages over traditional approaches, such as reducing the number of participating clients, utilizing smaller models, or solely training the classification head. Remarkably, the total communication overhead for $50$ rounds in the FedPEFT framework is less than that of a single round in a conventional FL setting.
  • In scenarios without communication limitations, FedPEFT can achieve server accuracies exceeding $95\%$ while only requiring less than $0.3\%$ of the parameters to be trainable.
  • The continuum of fine-tuning freedom, ranging from Full fine-tuning through FedPEFT-Bias, FedPEFT-Adapter, FedPEFT-Prompt, to Head-tuning, varies across different scenarios. Typically, FedPEFT-Bias and FedPEFT-Prompt emerge as the top performers in image and video processing tasks. Notably, FedPEFT-Prompt demonstrates superior adaptability when bridging larger domain gaps between pre-trained models and downstream tasks.
  • The effectiveness of FedPEFT is particularly pronounced in real-world applications characterized by stringent privacy requirements or data scarcity. Even under such conditions, FedPEFT-Bias and FedPEFT-Prompt maintain exceptional performance across image and video domains, surpassing other methods.
V Discussion
V-A Ablation Study
In the experiments above, we maintain a consistent backbone (ViT-B). However, since model size requirements of different federated systems may vary, we further investigate the use of ViT-S with FedPEFT and all baselines. Additionally, we study the effect of pre-training dataset size and paradigm on the final performance of all approaches. CIFAR-100 is used for the experiments.
Impact of Model Size. We replace ViT-B with a smaller version, ViT-S. As expected, most methods experience a slight performance drop. Nonetheless, FedPEFT methods maintain strong performance even with smaller models. Surprisingly, the performance of head-tuning is not damaged by reducing the model size but actually increases slightly. We interpret it as due to the fixed representation of head-tuning. Representations from a pre-trained model can have similar semantics whatever the model size, thus the performance of head-tuning is stable.
Impact of Pre-training Dataset. We also evaluate the performance when a backbone is trained from ImageNet-21K and ImageNet-1K [70]. Interestingly, the performance stays relatively consistent, with Head-tuning and FedPEFT-Adapter achieving higher accuracy with the ImageNet-1K models. This phenomenon indicates that the performance of fine-tuning is not always proportional to the data size, but overall FedPEFT performance is consistent with different scales of the pre-training dataset.
Impact of Pre-training Paradigm. Additionally, we evaluate one self-supervised pre-trained backbone: DINO [71] for the image domain. FedPEFT still achieves comparable results with full fine-tuning and outperforms Head-tuning, showing the robustness of FedPEFT.
Impact of Federated Learning Algorithm. Besides the widely-used FedAvg algorithm, FedPEFT focuses on the local training and communication stages, thus being orthogonal and compatible with other algorithms. We evaluate FedPEFT along with the baseline with FedProx [2] and MOON [21] in TABLE VIII and find the performance can be further improved when more advanced algorithms are integrated.
V-B Extension Capability
In this paper, we mainly focus on the computer vision fields with three base FedPEFT prototypes. We further show FedPEFT’s extension capability to be extended with other PEFT methods and for natural language processing tasks.
PEFT Methods. LoRA [29] and Prefix [34] are another popular backbone and input adaptation methods. As shown in TABLE IX, FedPEFT is compatible, and the performance of FedPEFT can be further improved with LoRA, while the communication costs can be further reduced with Prefix.
NLP task. We evaluate the performance of FedPEFT and baselines in text classification in the AG NEWS [72] dataset under the setting $M=64,N=8,\alpha=1$ with MiniBERT [73] as the pre-trained backbone in TABLE X. FedPEFT-Bias, with the highest fine-tuning freedom, still provides robust performance, which is consistent with our insights in CV. Further, a follow-up work of FedPEFT gives a study in NLP, further indicating the extension capability of FedPEFT [42].
TABLE X: Performance of FedPEFT on NLP task. # TP means the number of trainable parameters, and FedPEFT-B, -A, and -P mean FedPEFT prototypes with bias, adapter, and prompt.
Method Full Head FedPEFT-B FedPEFT-A FedPEFT-P
# TP 11M 1.0K 12.8K 17.7K 31.7K
Server Acc. 88.68 46.97 83.20 77.80 44.82
VI Conclusion
In this paper, we introduce FedPEFT, a new federated learning framework leveraging strong pre-trained models and massively reducing communication costs. We integrate three effective prototypes within the FedPEFT framework: Bias, Adapter, and Prompt. With a thorough empirical study, we then evaluate FedFEFT and other baselines in three key areas: communication, capability, and robustness. We find FedPEFT to be a promising approach for practical FL systems, capable of handling many of the harsh conditions in FL while alleviating the critical communication bottleneck. As a general framework, FedPEFT can also be leveraged with other PEFT methods and in application domains other than computer vision. We hope this work can inspire new perspectives in federated learning through the combined innovation of strong pre-trained models and parameter-efficient fine-tuning methodologies.
VII Acknowledgement
This work is partially supported by the NSF/Intel Partnership on MLWiNS under Grant No. 2003198.